OpenStudy (anonymous):

if Z = Cos x + i sin X Find \[\frac{z^2-1}{z^2 + 1}\]

OpenStudy (anonymous):

multiply num and denum by \(\overline{z}\)

OpenStudy (anonymous):

and note that \(|z|=1\)

OpenStudy (anonymous):

another hint : \(z\overline{z}=|z|^2=1\)

OpenStudy (anonymous):

come on yahoo!!!!

OpenStudy (anonymous):

\[\frac{(z^2 -1)(cosx-isinx)}{(z^2 + 1)(cosx-isinx)}\]

OpenStudy (anonymous):

lol....i was out for 10 min.. Sorry

OpenStudy (anonymous):

np

OpenStudy (anonymous):

\[\frac{z^2-1}{z^2 + 1}=\frac{z^2\overline{z}-\overline{z}}{z^2\overline{z} + \overline{z}}=\frac{zz\overline{z}-\overline{z}}{zz\overline{z} + \overline{z}}\]

OpenStudy (anonymous):

Oh....u told...like..this....)

OpenStudy (anonymous):

(z-z1)/(z+z1)

OpenStudy (anonymous):

yes and now simplify it

OpenStudy (anonymous):

i tanx

OpenStudy (anonymous):

Thxx...)))